During an adiabatic process pressure of gas is found to be proportional to the
cube of its temperature. The ratio of Cp/Cv is A. 2
B. 44319 C. 44257
D. 44289
1
Expert's answer
2021-10-03T17:13:53-0400

"\frac{C_p}{C_V}=\gamma,"

"T^{\gamma}p^{1-\gamma}=const,\implies"

"p\sim T^{\frac{\gamma}{\gamma-1}}," but

"p\sim T^3,\implies"

"\frac{\gamma}{\gamma-1}=3," where

"\gamma=\frac 32."